Smith is a weld inspector at a shipyard. He knows from keeping track of good and substandard welds that for the afternoon shift 4.5% of all welds done will be substandard. If Smith checks 250 of the 7500 welds completed in an afternoon shift, then what is the probability that he will find less than 15 substandard welds?
Answer #1

Let p be the proportion of the substandard welds and x be the number of subtandard welds.

Given : p = 0.045 , n = 250

1596248464420_blob.png = x /n = 15/250 = 0.06

According to sampling distribution of sample proportion 1596248520224_blob.png , The sample proportion approximately follows normal distribution with mean 1596248520244_blob.png = P  and standard deviation 1596248520224_blob.png = 1596248520248_blob.png   

Therefore 1596248520244_blob.png = 0.045

and 1596248520224_blob.png = \sqrt{\frac{0.045*(1-0.045)}{250}} = 0.0131

We are asked to find P( 1596248464420_blob.png < 0.06 )

= P\left ( \frac{\hat{p}-\mu _{\hat{p}}}{\sigma _{\hat{p}}} \leq\frac{0.06-0.045}{0.0131}\right )

= P( z ≤ 1.14 )

= 0.8729

The probability that he will find less than 15 substandard welds is 0.8729